Minutes of November 10, 2004 meeting
- Present:
- Megan Ingegno, Mary Brodman, Shirley Peck, Irina Loutchkina, Robert Krack, Edward Suarez, Matt Spaventa, Penny Weniger, Jeff Teichmann, Eugene McElroy, Melody Tomaszewicz, Martin Katz, Melanie Robinson, Krista Dandurand, Nicole Warren, Paul Young, Stacy DeMatteo, Andy Martinez, Brian Stubbs
- Excused:
- Rosely Riley
- Guest:
- Roger Smith
Agenda
- Information Sharing
- Water Emergency Supply Order
- Range Guide Template Project
- Internal User Project
- Next Meeting
INFORMATION SHARING
- Megan (Physics)- Howard Dess has visited to assess the collection in preparation for a
shift; Howard is also reviewing PERM reserve collection to see if items can be weeded to
create space to accommodate undergraduate reserves, which Physics currently does not have.
- Irina (Alex)- has begun processing reserve material for the Spring 2005 semester,
including the ordering of new books for the collection.
- Mary (Alex)- progress continues on several mini-shifts in the Stacks.
- Shirley (Alex)- since the introduction of the LST/ASS lists, there has been an
increase in the amount of claims returned and searching; has been receiving IN-PROCESS
books that are moldy, they’re suspected to be donations not new acquisitions; items that
are found are sent to Collection Services for review.
- Rob (Alex)- continues work on the mini-shifts; there are about 5 more sections to
complete by semester’s end.
- Eddie(LSM)- progress on weeding project continues; guides are being made for markers
placed as part of weeding project, to make locating items easier; X is being inventoried in
preparation for a transfer in the coming months.
- Matt (Douglass)- all items from the Folio Collection project have been sorted and are
now ready to be shelved; Alex CS will be contacted for a supply of DGLSS FOLIO labels.
- Penny (Douglass)- a shift proposal has been sent to Farideh for review; noticing an influx of damaged books being returned by patrons, particularly those who were billed for LST/ASS material.
- Melody (LSM-CSD)- continues to process items pulled as part of the LSM 2nd floor weeding
project; a statistics spreadsheet of the progress is being kept and updated monthly;
progress continues on the Marc Holdings project, most branches are 80%-100% finished with
phase 1 of the project.
- Marty (LSM)- continues to receive recataloged Gov Docs, along with the usual IN-PROCESS books.
- Melanie(Math)- working on the Marc Holdings project; slight increase in the amount of IN-PROCESS books being received.
- Krista (KLMR)- is now receiving and searching Stacks Check requests; working closely
with student employees on increasing their shelving and shelf reading skills.
- Nicole (Chang)- processing Cook Scholar honors theses, preparing them to be bound and
entering them into a database.
- Paul(Gov Docs)- a mini shift is occurring to accommodate the addition of the FCC
Records, at ALEX; recataloging of foreign/international documents continues at LSM.
- Stacy (Annex)- problems with water condensation continue, leading to water dripping on
collection from overhead beams; most recently 8 problem areas were found in one day;
continues to receive items from LSM weeding project.
WATER EMERGENCY SUPPLY ORDER
- The supply order for Water Emergency supplies was approved and orders have been placed
- A copy of the supplies ordered for individual units were distributed to those who
supplies were ordered for
RANGE GUIDE TEMPLATE
- A proposed template was distributed to the group for review.
- The group supports the idea of including numbers for individual ranges the guides but
questions how they could be displayed discreetly, without being confusing patrons; Andy
will consider several suggested options and share them with the group at a later date.
- The group proposes standardized color-coding of the guides according to specific
collections, but questions arose concerning the visibility of text on certain colors of
cardstock. Several options will be tested and considered. Decision of what color will
represent which collection will be decided at future meeting.
- Stacy will search for an Avery template (or similar product), that will allow for the
printing of 4 guides per sheet of cardstock.
INTERNAL USER PROJECT
- A list of internal users, currently not in use was distributed to the group. Items on the
list still have outstanding charges that need to be resolved.
- In the coming week, Andy will distribute reports indicating the owning library and the
library that last had contact with each book that is still checked out to these users.
Items are to be searched, discharged if found, and returned to their home locations. Any
items not found will be charged to the missing user of its home library.
NEXT MEETING
Wednesday December 8, 2004
2:30 p.m. at Kilmer or SMLR